8. Find the law of motion of a particle in the field U(x)=-Cx^4 ,if it is Total energy is zero ( C is positive constant)


1
Expert's answer
2021-06-22T09:47:30-0400

"F=U'=-4cx^3,"

"F=\frac Ux=-cx^3,"

"F=-5cx^3,"

"F=m\ddot x,"

"\ddot x=\frac Fm=-\frac{5cx^3}{m},"

"\dot x=-\frac{5cx^4}{4m},"

"x=-\frac{cx^5}{4m},"

"(-\frac{cx^5}{4m})^2-\omega^2\frac{5cx^3}{m}=0,"

"cx^7-80\omega^2 m=0."